# The SABRE South Technical Design Report Executive Summary

### v2

@@ 7.2 Optical calibration system
The veto detector system is monitored and calibrated by using optical light from a pulsed laser system: specifically the PMT timing, gain and resolution, optical properties of the LAB scintillator, and overall detector stability. The conceptual design of the optical calibration system uses optical fibres sheathed in Teflon tubes that carry picosecond laser pulses to designated emission positions within the veto vessel. Several configurations are considered. Figure 19 shows two of the considered configurations, where light is emitted from one and three diffusive light sources, respectively. Quartz fibres are chosen to transmit the light signal from the laser into the vessel as the material is chemically compatible with LAB. The optical fibres are sheathed with Teflon, to increase the reflectivity and improve signal isotropy, and fed into the veto vessel via one or more ConFlat 151.6 mm outer diameter nozzles (depending on the configuration).
The two optical calibration configurations were simulated with Geant4, to analyse the proportion of photoelectrons collected by the PMTs, and to examine the photon flight times, trajectories in the LAB, and their probability of survival. These configurations differ in the number of optical fibres and where they terminate. In both configurations, the fibres run along the edge of the vessel and terminate into a diffusive bulb held in place by a tripod. The diffuser bulbs are used to achieve an isotropic light source. For configuration 1, the bulb is located halfway between the bottom of the central enclosure and the bottom of the vessel, i.e. 65 cm directly below the centre of the vessel (which is the origin of the detector coordinate system). This position is optimised for the propagation of photons within the scintillator, as well as the number of photoelectrons detected by the PMTs. For configuration 2, the bulbs are arranged into an equilateral triangle centred on the origin. The distance from the origin to each bulb is 69 cm. The diffuser bulbs are solid spheres of PTFE with a diameter of about 1.5 cm, which is proven solution for LS experiments [30].
Figure 19: Left: configuration 1, calibration with a single diffusive light source located halfway between the bottom of the central enclosure and the vessel floor. Right: configuration 2, three evenly-spaced diffusive light sources for better coverage and redundancy within the veto vessel. The path of the optical fibre from the vessel ConFlat nozzle to the light emitter is shown in blue.
The nominal light source is an LDB-200 picosecond pulsed laser from Tama Electric, and can be remotely controlled. A wavelength 405 nm was initially considered due to the high quantum efficiency of PMTs at this wavelength. However, the photon absorption length at this wavelength is 0.3 m. Therefore a laser wavelength of 445 nm, where the absorption length is 25 m, is chosen. The PMT quantum efficiency is only 5% lower than at 405 nm.
[…]
For configuration 2, the proportion of detected photons is the same as the single source option, but there is a smaller difference in photons collected by each PMT across the detector. This configuration offers more timing calibration data by having multiple source positions. The nominal 445 nm laser source can output up to $2\times 10^{5}$ photons per pulse and reach a maximum pulse frequency of 100 MHz, allowing ample room for adjustment.
Both configurations are valid options for the optical calibration system. As configuration 1 is the easiest option to implement, we consider it the primary candidate.
@@ 8 Data Acquisition and slow control
A schematic layout of the Data Acquisition (DAQ) system is given in Fig. 20 as well as other related components that are discussed in more detail in later sections. Signals from every PMT are digitised by CAEN digitisers, and the resulting data are then transferred to a computer and written to disk. Each of the SABRE subsystems (the crystal, liquid scintillator, and muon detectors) is managed largely independently with three DAQ instances collecting data in parallel. For the crystal subsystem, a dual-readout approach is being considered in which two signals from each PMT are digitised (amplified and unamplified). A CAEN FPGA (Field Programmable Gate Array)-based board, running custom firmware, is used to coordinate triggering and provide synchronisation signals between the digitisers where required. A custom C++ program is responsible for managing the data collection from the digitisers. The system described here is operational in SUPL and used in precalibration (Fig. 21).
@@
Figure 20: Schematic layout of the SABRE DAQ, monitoring, and control systems.
@@ 8.1 Digitisers
LS veto V2730 32 $500\times 10^{6}$ 14
@@
The CAEN V1730 [33] is an 8-channel, 500 Ms/s digitiser with a 14-bit ADC and a 0.5/2 Vpp dynamic range. It is capable of continuous acquisition with zero dead time. Whenever the digitiser is triggered, the time of the event is recorded along with waveforms for each channel and some meta-data. To achieve an energy threshold of 1 keV, the trigger must be sensitive to single photoelectron signals. The noise rate at such a threshold demands the use of coincidence triggers across pairs of PMTs coupled to the same crystal and configured as channel pairs on the digitiser.
@@
Used to record signals generated in the LS veto, the 32-channel V7230 digitiser samples at 500 MS/s with a 14-bit ADC and a dynamic range of 4 Vpp [34]. The V2730 also has improved triggering options through custom FPGA firmware that this new generation of CAEN digitisers supports, allowing individual channel triggers, and correlated triggers between channels.
@@
The CAEN V1743 [28] is a 16-channel, 3.2 GS/s digitiser with a 12-bit ADC and a 2.5 Vpp dynamic range. It is used with the muon detector sub-system where the high-digitisation rate (and subsequent timing resolution) is useful for position reconstruction. However, due to its high digitisation rate, the board has a significant amount of dead time (125 $\mu$s for 1024 samples) per event. Since the expected event rate is $\sim$10 Hz and previous work with the muon detectors has acquired the full 1024 samples, this dead time is not expected to have a significant impact.
@@ 8.3 High voltage
The high voltage to the PMTs is supplied using a CAEN SY5527 [36] Universal Multichannel Power Supply System with A7435 [37] power supplies which can supply up to 3.5 kV and 3.5 mA per channel. Two 24-channel, positive polarity power supplies will be used for the crystal detectors and LS veto, whilst a single 24-channel negative polarity power supply will be used for the muon veto. These power supplies have a typical voltage ripple of < 7 mVpp, and voltage set resolution of 5 mV [37]. The mainframe has a built-in CPU controller to control and monitor each channel. The SY5527 mainframe has a built-in EPICS input-output controller (IOC).
@@ 8.5 Slow control
  • Vibration accelerometers: As SUPL is located in an active gold mine, there are occasions when vibrations from blasts or heavy vehicles elsewhere in the mine may be transmitted through to SUPL, inducing noise in the PMTs. Data on such vibrations are obtained via VSA001 single-axis accelerometers. Three of these are be mounted independently on the outside of the veto vessel, with each sensor oriented to provide x/y/z measurements. Analogue current output (0-10 mA) is proportional to motion (sensitive from 1 Hz to 6 kHz vibration frequency).
  • Radon: The level of radon in the SUPL environment will be monitored. During experiment operation, we are primarily concerned with radon in the vessel-shielding volume. This will be done via an AlphaGuard detector.
  • Oxygen: Oxygen sensors are placed within the shielding (external to the veto vessel) and within the nitrogen blanket inside the veto vessel (above the LAB level). There may also be oxygen monitoring within the fluid handling system and general SUPL environment. This is necessary as oxygen can dissolve into the LS and consequently reduce the light yield.
@@ 9 Shielding
Table 8 outlines the radiopurity of the shielding components. Samples of similar low-carbon aluminium-killed steel were screened by the Australian Nuclear Science and Technology Organisation (ANSTO) with a HPGe detector. No radioactive isotope was detected in the sample. This measurement was approximately ten times less sensitive than that used to detemine the measurements in Table 5, and therefore we consider conservative limits. The radioactivity of the HDPE layers is not expected to contribute significantly to background, so HDPE samples were not measured. Instead, limits based on measurements from XENON are used [40].
@@
Table 8: Radioactivity levels of the shielding steel and HDPE components, based on steel radioactivity measured by ANSTO and HDPE values from XENON [40].
@@ 10 Expected sensitivity of SABRE South
The key performance parameters are driven by the fact that, for NaI(Tl) crystals, a 1 keV deposition produces 30-45 photons, translating to 10-15 photoelectrons detected across the two PMTs in each module. The SABRE South region of interest therefore has a series of distinct single-photon pulses distributed over several hundreds of nanoseconds. All currently operating NaI(Tl) experiments have background models that poorly reproduce the data at energies below 3 keV [41, 2]. Part of the difficulty in modelling the background in this low-energy region is due to the challenges of separating PMT noise, caused by spontaneous thermionic emission of electrons, from a genuine scintillation event of interest.
@@ 11 Summary and outlook
This document presents the executive summary of the technical design of the SABRE South detector, a NaI(Tl)-based detector that aims to test the long-standing DAMA/LIBRA annual modulation result. Using the same target material as the DAMA/LIBRA detector, the SABRE South NaI(Tl) crystals have an ultra-low background, owing to growth from Merck Astrograde NaI powder (which has potassium contamination below 10 ppb). Crystal growth is ongoing through collaborations with SICCAS (China) and RMD (USA). Each crystal is coupled to two PMTs and enclosed in an oxygen-free, high-thermal-conductivity copper enclosure that is submerged in the SABRE South vessel (suspended by copper conduits). The vessel itself contains 11,600 L of liquid scintillator – a mixture of LAB doped with approximately 3.5 g/L of PPO and 15 mg/L of bis-MSB. The vessel is instrumented with a baseline of 18 PMTs, and as many as 32. By including a LS veto detector, and submerging the crystal detector modules within in it, radioactive decays from within the crystals that may emit high-energy gamma rays can be tagged — for example, 40K decays, which form a key impurity in the NaI(Tl) crystals. The LS veto detector can also veto external background, with the possibility to employ reconstruction methods to disentangle background sources. In addition to the LS veto, a 9.6 m2 muon veto will be placed above the detector and shielding to veto incoming cosmogenic muons. This muon veto is also capable of position reconstruction of tagged muons along the length of the panels, and can be used in coincidence with the LS veto.
@@ Acknowledgements
The SABRE South program is supported by the Australian Government through the Australian Research Council (Grants: CE200100008, LE190100196, LE170100162, LE160100080, DP190103123, DP170101675, LP150100705). This research was partially supported by Australian Government Research Training Program Scholarships, and Melbourne Research Scholarships. This research was supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative. We thank the SABRE North collaboration for their contributions to the SABRE South experiment design, engineering, NaI(Tl) crystal development, background screening, and simulation framework. Work and equipment installation in the Stawell Underground Physics Laboratory was supported by the SUPL Board.
@@
RCS Radioactive calibration system

### v3

@@ 7.2 Optical calibration system
The veto detector system is monitored and calibrated using optical light from a pulsed laser system. This calibration involves assessing the PMT timing, gain, resolution, optical properties of the LAB scintillator, and overall detector stability. The conceptual design of the optical calibration system utilises optical fibres, which are sheathed in Teflon tubes, to transport picosecond laser pulses to specific emission positions within the veto vessel.
Two diffusive bulbs (or laser balls) are attached to steel calibration tubes, which are then connected to ConFlat 151.6 mm outer diameter flanges on the veto vessel. Quartz optical fibres are chosen to carry signals from the laser light source, as the material is chemically compatible with LAB. The optical fibres are sheathed with PTFE to increase reflectivity and improve signal isotropy. The laser balls are nominally positioned at two different vertical and horizontal positions within the vessel. This configuration allows for studies of single photoelectron signals with each PMT. They are made of PTFE with radii of 25 mm and are a proven solution for liquid scintillator experiments [55]. The fibres are attached to the laser balls using connectors made of LAB-compatible polyether-ether-ketone resin (PEEK) [55]. Studies on shadowing effects caused by the connector have shown that this effect is negligible. The stainless steel tubes provide stability to the system.
The laser is located outside the detector, with an optical splitter dividing the beam into three fibres: two fibres are inserted into the vessel, and one connects to a reference sensor system. The nominal light source is an LDB-200 picosecond pulsed laser from Tama Electric, which can be remotely controlled. A wavelength of 405 nm was initially considered due to the high QE of the PMTs at this wavelength. However, the photon absorption length at this wavelength is 0.3 m. As a result, a laser wavelength of 445 nm, where the absorption length is 26 m, is chosen. The PMT QE is only 5% lower than at 405 nm.
The optical calibration system is simulated with Geant4 to ensure that it is suitable for purpose. The simulation has been used to determine the number of photoelectrons collected by each PMT, and to examine the expected photon flight times, trajectories in the LAB, and their probability of survival.
@@ 8 Data Acquisition and slow control
A schematic layout of the Data Acquisition (DAQ) system is given in Fig. 21 as well as other related components that are discussed in more detail in later sections. Signals from every PMT are digitised by CAEN digitisers, and the resulting data are then transferred to a computer and written to disk. Each of the SABRE subsystems (the crystal, liquid scintillator, and muon detectors) is managed largely independently with three DAQ instances collecting data in parallel. For the crystal subsystem, a dual-readout approach is being considered in which two signals from each PMT are digitised (amplified and unamplified). A CAEN FPGA (Field Programmable Gate Array)-based board, running custom firmware, is used to coordinate triggering and provide synchronisation signals between the digitisers where required. A custom C++ program is responsible for managing the data collection from the digitisers. The system described here is operational in SUPL and used in pre-calibration (Fig. 22).
@@
Figure 21: Schematic layout of the SABRE DAQ, monitoring, and control systems. Note that this is for the nominal crystal readout not including preamplifiers. Dual-readout with preamplifiers on the crystals would require 2 additional V1730 digitisers.
@@ 8.1 Digitisers
Liquid scintillator veto V2730 32 $500\times 10^{6}$ 14
@@
The CAEN V1730 [58] is an 8-channel, 500 MS/s digitiser with a 14-bit ADC and a 0.5 or 2 Vpp (peak-to-peak) dynamic range. It is capable of continuous acquisition with zero dead time. Whenever the digitiser is triggered, the time of the event is recorded along with waveforms for each channel and some meta-data. To achieve an energy threshold of 1 keV, the trigger must be sensitive to single photoelectron signals. The noise rate at such a threshold demands the use of coincidence triggers across pairs of PMTs coupled to the same crystal and configured as channel pairs on the digitiser.
@@
Used to record signals and waveforms generated in the liquid scintillator veto, the 32-channel V7230 digitiser samples at 500 MS/s with a 14-bit ADC and a dynamic range of 4 Vpp [59]. The V2730 also has improved triggering options through custom FPGA firmware that this new generation of CAEN digitisers supports, allowing individual channel triggers, and correlated triggers between channels.
@@
The CAEN V1743 [50] is a 16-channel, 3.2 GS/s digitiser with a 12-bit ADC and a 2.5 Vpp dynamic range. It is used with the muon detector sub-system where the high-digitisation rate (and subsequent timing resolution) is useful for position reconstruction. Due to its high digitisation rate, the board has a significant amount of dead time (125 $\mu$s for 1024 samples) per event. Since the expected event rate is $\sim$10 Hz (estimated based on measurements, inclusive of noise, of the muon detectors installed in SUPL) the dead time is not expected to have a significant impact on efficiency.
@@ 8.3 High voltage
The high voltage to the PMTs is supplied using a CAEN SY5527 [61] Universal Multichannel Power Supply System with A7435 [62] power supplies which can supply up to 3.5 kV and 3.5 mA per channel. Two 24-channel, positive polarity power supplies will be used for the crystal detectors and liquid scintillator veto, whilst a single 24-channel negative polarity power supply will be used for the muon veto. These power supplies have a typical voltage ripple of < 7 mVpp, and voltage set resolution of 5 mV [62]. The mainframe has a built-in CPU controller to control and monitor each channel. The SY5527 mainframe has a built-in EPICS input-output controller (IOC).
@@ 8.5 Slow control
  • Vibration accelerometers: As SUPL is located in an active gold mine, there are occasions when vibrations from blasts or heavy vehicles elsewhere in the mine may be transmitted through to SUPL, inducing noise in the PMTs. Data on such vibrations are obtained via VSA001 single-axis accelerometers. Three of these are be mounted independently on the outside of the veto vessel, with each sensor oriented to provide x/y/z measurements. Analogue current output (0–10 mA) is proportional to motion (sensitive from 1 Hz to 6 kHz vibration frequency).
  • Radon: The level of radon in the SUPL environment will be monitored. During experiment operation, we are primarily concerned with radon in the vessel-shielding volume. This will be done with an AlphaGuard DF2000 detector with a measurement range of 2–2,000,000 Bq/m3 and sensitivity of 5 counts per minute at 100 Bq/m3.
  • Oxygen: Oxygen sensors are placed within the shielding (external to the veto vessel) and within the nitrogen blanket inside the veto vessel (above the LAB level). There may also be oxygen monitoring within the fluid handling system and general SUPL environment. This is necessary as oxygen can dissolve into the liquid scintillator and consequently reduce the light yield.
The temperature is controlled via a whole-laboratory solution. The laboratory is cooled to approximately 23 C. Humidity and radon within and around the detector elements is controlled via constant flushing of HPN. HPN is constantly flushed through the crystal modules, the volume between the shielding and the vessel, and through a blanket on top of the liquid scintillator within the vessel.
@@ 9 Shielding
Table 8 outlines the radiopurity of the shielding components. Samples of similar low-carbon aluminium-killed steel were screened by the Australian Nuclear Science and Technology Organisation (ANSTO) with a HPGe detector. No radioactive isotope was detected in the sample. This measurement was approximately ten times less sensitive than the measurements in Table 5, and therefore we consider conservative limits. The radioactivity of the HDPE layers is not expected to contribute significantly to background, so HDPE samples were not measured. Instead, limits based on measurements from XENON are used [65].
@@
Table 8: Radioactivity levels of the shielding steel and HDPE components [11], based on steel radioactivity measured by ANSTO and HDPE values from XENON [65].
@@ 10 Expected sensitivity of SABRE South
The key performance parameters are driven by the fact that, for NaI(Tl) crystals, a 1 keV deposition produces 30–45 photons, translating to 10–15 photoelectrons detected across the two PMTs in each module (based on a calculation using the light yield quoted in Table 9, and a PMT QE of 25–30%). The SABRE South region of interest therefore has a series of distinct single-photon pulses distributed over several hundreds of nanoseconds. All currently operating NaI(Tl) experiments have background models that poorly reproduce the data at energies below 3 keV [21, 15, 32]. Part of the difficulty in modelling the background in this low-energy region is due to the challenges of separating PMT noise, caused by spontaneous thermionic emission of electrons, from a genuine scintillation event of interest. Contributions from Cherenkov emissions in the glass of the crystal PMTs, and crystal afterglow following high intensity signals, are also being examined.
@@ 11 Summary and outlook
This document presents the executive summary of the technical design of the SABRE South detector, a NaI(Tl)-based detector that aims to test the long-standing DAMA/LIBRA annual modulation result. Using the same target material as the DAMA/LIBRA detector, the SABRE South NaI(Tl) crystals have an ultra-low background, owing to growth from Merck Astrograde NaI powder (which has potassium contamination below 10 ppb). SABRE South is also uniquely positioned in the Southern Hemisphere, allowing for potential disentanglement of seasonal effects. Crystal growth is ongoing through collaborations with SICCAS (China) and RMD (USA). Each crystal is coupled to two PMTs and enclosed in an oxygen-free, high-thermal-conductivity copper enclosure that is submerged in the SABRE South vessel (suspended by copper conduits). The vessel itself contains 11,600 L of liquid scintillator — a mixture of LAB doped with approximately 3.5 g/L of PPO and 15 mg/L of bis-MSB. The vessel is instrumented with a baseline of 18 PMTs, and as many as 32. By including a liquid scintillator veto detector, and submerging the crystal detector modules within it, radioactive decays from within the crystals that may emit high-energy gamma rays can be tagged — for example, 40K decays, which form a key impurity in the NaI(Tl) crystals. The liquid scintillator veto detector can also veto external background, with the possibility to employ reconstruction methods to disentangle background sources. In addition to the liquid scintillator veto, a 9.6 m2 muon veto will be placed above the detector and shielding to veto incoming cosmogenic muons. This muon veto is also capable of position reconstruction of tagged muons along the length of the panels, and can be used in coincidence with the liquid scintillator veto.
@@ Acknowledgements
The SABRE South program is supported by the Australian Government through the Australian Research Council (Grants: CE200100008, LE190100196, LE170100162, LE160100080, DP190103123, DP170101675, LP150100705). This research was partially supported by Australian Government Research Training Program Scholarships, and Melbourne Research Scholarships. This research was supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative. We thank the SABRE North collaboration for their contributions to the SABRE South experiment design, in particular the crystal enclosure, the crystal insertion system and the gas handling system; and to the NaI(Tl) crystal development, background screening, and simulation framework. Work and equipment installation in the Stawell Underground Physics Laboratory is supported by the SUPL Board. We thank the Institute of High Energy Physics (IHEP), China, for their help in the procurement of the liquid scintillator, and for graciously providing us with the additional PMTs from the decommissioned Daya Bay experiment. For work on the crystal production, we thank RMD, Boston, and SICCAS, Shanghai, as well as Mellen (for the zone-refining process). For the production of the NaI-33 crystal, we thank Frank Calaprice and Jay Benziger from Princeton, alongside Seastar for providing the ICP-MS measurements, and Sandfire for providing the crucible in which the crystal was grown.
@@
RCS Radioisotope calibration system